OpenStudy (anonymous):

if Re(z) = 0 then is z a purely real number or imaginary number

OpenStudy (diyadiya):

i hope its Imaginary number

OpenStudy (anonymous):

yes u r right but how diya ?

OpenStudy (diyadiya):

Re(z) = Real part of z let z be a+ib Re(z)= a im(z)=b Re(z)=0 therefore a=0 z=0+ib =ib and its an imaginary number

OpenStudy (diyadiya):

Re(z) = Real part of z Im(z)=Imaginary part of z

OpenStudy (diyadiya):

did you understand?

OpenStudy (anonymous):

yes diya thanks a lot

OpenStudy (diyadiya):

welcome:)

OpenStudy (anonymous):

You forgot something, what if z=0? You have to separate in two cases.

OpenStudy (diyadiya):

Well ,isn't it given Re(z)=0 ?

OpenStudy (anonymous):

When Re9Z0=0 the result is imaginary, on the other side if Im9z)=0, it is real number.Got it ?

OpenStudy (anonymous):

1348 is right. In the case Im(z)=0 then z=0+i0=0, which is in R.

OpenStudy (anonymous):

SRY Re(z)

OpenStudy (anonymous):

Though the implication is correct, since z is either real or complex, exclusively.

OpenStudy (anonymous):

if Im(z) = 0, then z is a real number. On the other hand, if RE(z)=0 it is an imaginary number.

OpenStudy (anonymous):

Two cases: \[1. z \in \mathbb{C}^*\] \[Re(z)=0 \iff z=re^{i \theta}, \theta \in ]-\pi ; \pi [ \setminus \left\{ 0 \right\}, r \in \mathbb{R}^* \\\] \[\iff z \in i\mathbb{R} \] BUT, case two... \[z=0 \in \mathbb{R} \implies Re(z)=0\] So no, your implication isn't direct. I can have a real number z with Re(z)=0, that number's called 0!

OpenStudy (anonymous):

\[z \in i\mathbb{R} \implies Re(z) =0 \] is true BUT, here's the difference... \[Re(z)=0 \implies z \in i\mathbb{R} \] is false, because 0 is real and Re(0)=0

OpenStudy (anonymous):

its imaginary number ,,but overall it is a complex number ,, as u know every number real or imaginary can be expressed in complex form so it is always suitable to say such numbers complex numbers ,,,but if u want to be specific it is imaginary number .

OpenStudy (anonymous):

why can i give only 1 medal for u all thanks a lot

OpenStudy (diyadiya):

OS Changed :)